OpenStudy (anonymous):

Help me resolve this equation? h (t) = -0.6(t-1.5)^2 + 1.35 We are looking for the value of t. h = 1

OpenStudy (anonymous):

are you trying to solve \[-0.6(t-1.5)^2 + 1.35 =1\]?

OpenStudy (anonymous):

No, I am trying to find the value of t.

OpenStudy (anonymous):

subtract 1.35 from both sides to get \[h (t) = -0.6(t-1.5)^2 =-.35 \]

OpenStudy (anonymous):

yes i see, we are looking for t.

OpenStudy (anonymous):

then divide by -.06 to get \[(t-1.5)^2=\frac{7}{12}\]

OpenStudy (anonymous):

so \[t-1.5=\pm\sqrt{\frac{7}{12}}\]

OpenStudy (anonymous):

and finally \[t=1.5\pm\sqrt{\frac{7}{12}}\]

OpenStudy (anonymous):

Thank you! That helped a lot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I have one more question. When you found 1.35, and put it on the other side to get -.35, where did the 1 go? Thanks again